OpenStudy (emmy1):

PLEASE HELP. END OF YEAR AND NEED GOOD GRADES! WILL FAN AND MEDAL Solve the system of equations by using the substitution method. Check each solution. Show your work. 3x + 2y = 6 x + 4y = -8

OpenStudy (cwrw238):

from second equation x = -8 - 4y now substitute for x in the first equation 3(-8-4y) + 2y = 6 -24 - 12y + 2y = 6 -10y = 30 can you continue? find value of y from the last line then plug this into the second equation to find x then plug the values of x and y into both the original equations to see if they fit.

OpenStudy (anonymous):

Mathy stuff: http://prntscr.com/3plm0q

OpenStudy (emmy1):

Thanks!!!